7. Study Questions # 14. Ch 2. The following table shows hypothetical export price indexes and import price indexes (2000 -100) for Japan, Canada, and Ireland. Compute the commodity terms of trade for each country for the period 2000-2016. Export Price Index Import Price Index Country 2000 2016 2000 2016 170 175 190 lapan Canada Treland 100 100 100 180 175 167 100 100 100 Which of the following accurately state how a countrys terms of trade improved, worsened, or showed no change? Check all that apply. Irelands commodity terms of trade remained constant at 100. Irelands commodity terms of trade worsened to 88. Japans commodity terms of trade worsened to 88. Japans commodity terms of trade improved to 106.
Answer #1

Solution:

A country's terms of trade = (Country's export price index/country's import price index)*100

In the year 2000, for all three countries: Japan, Canada, Ireland = (100/100)*100 = 100

In the year 2016, terms of trade (TOT) for

JapanTOT = (180/170)*100 = 105.88

CanadaTOT = (175/175)*100 = 100

IrelandTOT = (167/190)*100 = 87.89

Clearly, Japan's TOT has improved to 106 (approx), Canada's TOT remained constant at 100, and Ireland's TOT has worsened to 88 (approx). So, among the given options, option (b) Ireland's commodity TOT worsened to 88, and (d) Japan's commodity TOT improved to 106. Thus, the options to be checked are options (b) and (d).